Suppose the short-run production function is q = 8L0.5. If the marginal cost of producing the 20th unit is $12, what is the wage per unit of labor?
Answer #1

The short-run production function is q = 8L^(0.5)

The marginal product of the labor is MPL = dq/dL = 8*(0.5)L^(0.5-1)

=4L^(-0.5)

At L = 20, the marginal product of the labor is MPL = 4L^(-0.5)

=4*(20)^(-0.5)

=0.894427191

MC(L=20) = $12

Hence, the wage per unit of labor = W = MC*MPL = $12*0.894427191 = $10.73312629

Thus, the wage per unit of labor is $10.73312629